OpenStudy (anonymous):

x-4/3 + 11/6 = x+1/2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

x-4/3 + 11/6 = x+1/2 x-8/6 + 11/6 = x+3/6 3/6 = 3/6 1/2 =1/2